# Expert-elicitation method for non-parametric joint priors using normalizing flows

### v1

@@ 4.3.3 Scenario 3: Correlated Normal Priors
In this final scenario, we introduced a specific dependency structure among the regression coefficients, while all other aspects remained identical to Scenario 1 (M2). The results of the sensitivity analysis, provided in Appendix A.4, indicate that the current set of “elicited” statistics would not provide sufficient information to learn the dependency structure. To address this, we included the pairwise correlations between model parameters as input to the learning algorithm. As a result, the correlation should be accurately learned, given the full information provided.
Fitting Stage
The convergence check indicated successful convergence, with results provided in Appendix A.4. The corresponding match between the model-implied and true “elicited” statistics is shown in Figure 12. We observe a close match for the quantiles of the predictive distribution of each group, while, as noted already in Scenario 2, the quantiles for $R^{2}$ remain more challenging to learn accurately. However, in this scenario, the overestimation of $R^{2}$ is less pronounced, occurring in only one replication, while all other replications appear to learn this statistic accurately. The correlation plot (shown on the bottom of Figure 12) also demonstrates good learning of the dependency information. Therefore, with the provided correlation information, the learning algorithm is able to learn a correlated joint prior.
@@
Figure 12: Normal model / Scenario 3 (M4) — Learned elicited statistics for all 30 replications: Upper plots show the learned quantiles (y-axis) vs. the true quantiles (x-axis) of each replication for the four target quantities $y\mid\text{gr}_{i}$ with $i=1,2,3$ and $R^{2}$. The diagonal line serves as a reference, where points lying on the line indicate a perfect match between learned and true quantiles. The lower plot displays the learned (orange) vs. true (red) correlations between the model parameters. All learned elicited statistics match the ground truth, with some variation observed between replications.
Evaluation Stage
In Figure 13b, the learned marginal priors for each replication are shown, with the computed weights used for model averaging displayed in Figure 13a. As expected, the learned marginal priors for the intercept parameter, $\beta_{0}$, show little variation, while there is substantial variation in the learned marginal priors for the other model parameters, whereby the variation for $\beta_{1}$ is less pronounced compared to $\beta_{2}$. As previously noted through the elicited statistics, there is again one replication that learns a marginal prior for $\sigma$ that places significantly more probability mass on values near zero compared to the learned marginal priors of the other replications. These results again suggest some multi-modality in the learned priors, which could be useful for constraining the search space if certain sets of priors are deemed unrealistic by the domain expert.
@@
Figure 13: Normal model / Scenario 3 (M4) — Learned marginal priors for all 30 replications: (a) Computed weights per replication used for model averaging, reflecting the relative performance of each model based on the final total loss. The highest and lowest weights are highlighted in green and orange, respectively. (b) The learned marginal priors for each replication are shown in grey, with the dotted black line representing the ground truth and the dashed black line the model average. The marginals for the seeds with the highest and lowest weights are highlighted in the respective colors.
@@ 5 Discussion & Outlook
In this paper, we introduce an elicitation method for learning non-parametric joint priors building on and extending our simulation-based framework [6]. This work demonstrates that our framework is capable of supporting the development of various elicitation methods, in addition to the previously introduced method for learning independent parametric priors. To learn non-parametric joint priors, we utilize normalizing flows [44] to learn flexible joint priors for the model parameters. Through four simulation studies, we demonstrated the successful learning of such priors informed by expert knowledge. Although we employ normalizing flows in this work, other generative models, such as flow matching [45] or diffusion models [9] are also compatible with our framework and can be readily used as a drop-in replacement.
We view the ability to use the same framework for implementing various types of elicitation methods advantageous. By constructing these methods from shared modules (building blocks), our approach not only facilitates direct comparison between elicitation methods but also supports the establishment of benchmarks within the field. Furthermore, a unified framework facilitates the development of a standardized set of diagnostics and analysis tools, addressing a gap that currently exists in this area [50]. These tools can be employed during the elicitation process to specify target quantities, choose appropriate elicitation techniques, and evaluate the learned prior distributions.
@@
Within our framework, expert knowledge can be specified in various ways—ranging from querying model parameters to observable quantities or any derived statistic—tailored to the specific problem and the expertise of the domain expert being queried. This flexibility is achieved through the inclusion of a module that defines both the quantities to be elicited from the expert (cf. target quantities) and the techniques used to query these quantities (cf. elicitation techniques). This design choice provides users with maximum flexibility in specifying the set of quantities to be elicited from the domain expert. However, it also places the responsibility on the user to define a valid and appropriate set of quantities to query from the expert [68]. We would argue that this is in general a desirable property of a method, as it avoids imposing specific modeling choices on the user. Instead of requiring that the problem is framed in a predefined way, the method allows users to adapt it to their specific needs [64, 68, 66]. However, in practice, this flexibility is only effective if good recommendations, guidelines, and diagnostics are available to help users define an appropriate set of quantities to query from the expert [29].
@@
A significant challenge in this context is assessing whether the chosen set of statistics is sufficiently informative for the corresponding model. The goal is not necessarily to identify a unique prior that fully defines the generative model. Instead, the aim is to determine a set of equivalently plausible priors, which is theoretically justified since a domain expert is unlikely to have a single “true” prior in mind [73]. This introduces at least two additional requirements for an elicitation method. First, it should include an efficient implementation that allows the training algorithm to be executed multiple times with different seeds, enabling the assessment of variability in the learned prior distributions. Second, the elicitation method should ideally offer additional diagnostics or advanced techniques, such as model selection or model averaging, to assist the user [23]. In this work, we propose a model averaging method in which the weights calculated for averaging can also serve as diagnostics, helping to identify potential outliers within the set of learned priors. Identifying these edge cases could provide a valuable starting point for evaluating the plausibility of the learned priors in collaboration with a domain expert.
@@
Another important direction for future work is the development of a technique that can learn the dependency structure of a joint prior while simultaneously querying the expert about interpretable quantities. In this paper, we demonstrated that our method can learn a joint prior with the corresponding dependency structure when informed with the pairwise model parameter correlation. However, for practical applicability, the goal is to build on these results and develop a method that infers the correlation indirectly from the expert by asking about interpretable quantities. Recently, Mikkola et al [51] proposed a method that uses pairwise comparisons or ranking techniques to learn a multivariate prior density, which could serve as an interesting avenue for future research.
Other important design choices within our framework that require appropriate default settings include the specification of the discrepancy measure and the selection of the optimization approach, among other aspects. So far, in all simulation studies, we have used the maximum mean discrepancy [MMD, 36] to evaluate the discrepancy between the model-implied and expert-elicited statistics. The only exception is the correlation information, for which we employed an L2 loss. One key advantage of the MMD is its flexibility in handling a wide range of statistics, from simple point estimates to complex histogram data. However, a notable limitation of the MMD is its high (i.e., quadratic) computational cost. In contrast, the L2 loss is computationally efficient but lacks the capability to process histogram data. Despite the higher computational cost, the MMD consistently showed very good performance and robustness across all simulation studies so far. Therefore, it appears to be a solid choice as a default discrepancy measure. Future work should study the performance of different loss functions depending on different elicited statistics, with the goal to develop user guidelines for selecting an appropriate loss function.
@@ Acknowledgements
We thank Dmytro Perepolkin and Luna Fazio for their valuable comments and suggestions that greatly improved the manuscript. All authors gratefully acknowledge the computing time provided on the Linux HPC cluster at Technical University Dortmund (LiDO3), partially funded in the course of the Large-Scale Equipment Initiative by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) as project 271512359.
@@
The simulation results are available at https://osf.io/xrzh6/.
@@
Not applicable.
@@
[…]
@@
Conceptualization: PB, SR, FB, Methodology: PB, FB, Software: FB, Writing (Original Draft): FB, Writing (Review & Editing): SR, PB, Visualization: FB, Supervision: PB, Funding acquisition: PB, SR
@@ A.1 Training time
The average training time across the 30 replications was
  • Binomial model: 24 minutes and 10 seconds ($\pm$ 33 seconds).
  • Independent normal scenario: 66 minutes and 50 seconds ($\pm$ 5 min: 53 secs).
  • Skewed normal scenario: 74 minutes and 34 seconds ($\pm$ 9 min: 16 secs).
  • Correlated normal scenario: 73 minutes and 18 seconds ($\pm$ 1 min: 59 secs).
@@ A.2 Independent normal scenario
Figure 14: Normal model / Scenario 1 (M2) — Sensitivity Analysis: Rows represent hyperparameters of each model parameter. Columns represent elicited statistics: five quantiles for $y\mid\text{gr}_{i}$ with $i=1,2,3$ and for $R^{2}$. Quantiles are depicted in different colors. In each row, the corresponding hyperparameter, is varied across the range shown on the x-axis, while all other hyperparameters are held constant at their true values, indicated by the red vertical line. The horizontal gray line is included as a reference to distinguish between the different model parameters.
Figure 15: Normal model / Scenario 1 (M2) — Preliminary Convergence Check for all 30 replications: The upper plot displays the absolute slope of the total loss trajectory (scaled by a factor of 100) over the last 100 epochs on the y-axis, with each replication shown along the x-axis. The replications (i.e., seeds) are arranged in ascending order based on the magnitude of the absolute slope. The best-performing model, with an almost zero slope, is seed 13, while the worst-performing models, indicated by negative slopes, are seeds 11, 2, 12, and 28. The lower plot illustrates the loss trajectories of the best- and worst-performing seeds over the last 200 epochs. The loss updates for the final 100 epochs are highlighted in green for the best model and in blue for the worst models, with a red line segment indicating the respective slope.
Figure 16: Normal model / Scenario 1 (M2) — Visual convergence check for seed=28: (a) Updating trajectory across epochs for various quantities of interest: the total loss (upper left), individual loss components (upper right), and the mean and standard deviation of the marginal priors (lower left and lower right, respectively). (b) Learned joint prior with marginal distributions along the diagonal and correlation information in the off-diagonal.
@@ A.3 Skewed normal scenario
Figure 17: Normal model / Scenario 2 (M3) — Sensitivity Analysis: Rows represent hyperparameters of each model parameter. Columns represent elicited statistics: five quantiles for $y\mid\text{gr}_{i}$ with $i=1,2,3$ and for $R^{2}$. Quantiles are depicted in different colors. In each row, the corresponding hyperparameter, is varied across the range shown on the x-axis, while all other hyperparameters are held constant at their true values, indicated by the red vertical line. The horizontal gray line is included as a reference to distinguish between the different model parameters.
Figure 18: Normal model / Scenario 2 (M3) — Preliminary Convergence Check for all 30 replications: The upper plot displays the absolute slope of the total loss trajectory (scaled by a factor of 100) over the last 100 epochs on the y-axis, with each replication shown along the x-axis. The replications (i.e., seeds) are arranged in ascending order based on the magnitude of the absolute slope. The best-performing model, with an almost zero slope, is seed 21, while the worst-performing models, indicated by negative slopes, are seeds 15, 7, 8, and 27. The lower plot illustrates the loss trajectories of the best- and worst-performing seeds over the last 200 epochs. The loss updates for the final 100 epochs are highlighted in green for the best model and in blue for the worst models, with a red line segment indicating the respective slope.
Figure 19: Normal model / Scenario 2 (M3) — Visual convergence check for seed=27: (a) Updating trajectory across epochs for various quantities of interest: the total loss (upper left), individual loss components (upper right), and the mean and standard deviation of the marginal priors (lower left and lower right, respectively). (b) Learned joint prior with marginal distributions along the diagonal and correlation information in the off-diagonal.
@@ A.4 Correlated normal scenario
Figure 20: Normal model / Scenario 3 (M4) — Sensitivity Analysis: Rows represent hyperparameters of each model parameter. Columns represent elicited statistics: five quantiles for $y\mid\text{gr}_{i}$ with $i=1,2,3$ and for $R^{2}$. Quantiles are depicted in different colors. In each row, the corresponding hyperparameter, is varied across the range shown on the x-axis, while all other hyperparameters are held constant at their true values, indicated by the red vertical line. The horizontal gray line is included as a reference to distinguish between the different model parameters.
Figure 21: Normal model / Scenario 3 (M4) — Preliminary Convergence Check for all 30 replications: The upper plot displays the absolute slope of the total loss trajectory (scaled by a factor of 100) over the last 100 epochs on the y-axis, with each replication shown along the x-axis. The replications (i.e., seeds) are arranged in ascending order based on the magnitude of the absolute slope. The best-performing model, with an almost zero slope, is seed 19, while the worst-performing models, indicated by negative slopes, are seeds 14, 5, 27, and 6. The lower plot illustrates the loss trajectories of the best- and worst-performing seeds over the last 200 epochs. The loss updates for the final 100 epochs are highlighted in green for the best model and in blue for the worst models, with a red line segment indicating the respective slope.
Figure 22: Normal model / Scenario 3 (M4) — Visual convergence check for seed=26: (a) Updating trajectory across epochs for various quantities of interest: the total loss (upper left), individual loss components (upper right), and the mean and standard deviation of the marginal priors (lower left and lower right, respectively). (b) Learned joint prior with marginal distributions along the diagonal and correlation information in the off-diagonal.

### v2

@@ 4.3.3 Scenario 3: Correlated Normal Priors
In this final scenario, we introduced a dependency structure among the regression coefficients, while all other aspects remained identical to Scenario 1 (M1). The results of the sensitivity analysis and the convergence checks are provided in Appendix A.4 and show successful convergence. The model-implied and true “elicited” statistics are shown in Figure 12 and indicate a close match for all elicited statistics.
@@
Figure 12: Normal model / Scenario 3 (M4) — Learned elicited statistics for all 30 replications: First four plots show the learned quantiles (y-axis) vs. the true quantiles (x-axis) of each replication for the four target quantities $y\mid\text{gr}_{i}$ with $i=1,2,3$ and $R^{2}$. The diagonal line serves as a reference, where points lying on the line indicate a perfect match between learned and true quantiles. The rightmost plot displays the learned (black) vs. true (red) correlations between the model parameters.
In Figure 13, the final learned marginal priors for each replication are shown alongside the results from prior averaging. Results show substantial variation in the learned priors for $\beta_{2}$, whereas the priors for all other parameters show relatively high consistency across replications. Since the focus of this case study is specifically on the correlation structure of the regression coefficients, it is informative to also examine the joint prior distribution, which is shown in Figure 14. The joint priors from different replications are shown in different colors. Although there is clear variation in the learned joint priors, the method is able to capture the underlying correlational pattern. Although this result is not particularly surprising, as we provided the method with the exact correlation information between the model parameters, it nonetheless indicates that the method is capable of learning the correlational structure of a joint prior when provided with the relevant information. Therefore, future work is needed to develop elicitation methods that enable the method to obtain relevant information by querying domain experts indirectly.
@@
Figure 13: Normal model / Scenario 3 (M4) — Learned marginal priors for all 30 replications: Upper plot: Computed weights per replication used for model averaging, reflecting the relative performance of each model based on the final total loss. Lower plot: The learned marginal priors for each replication (black) and the prior average (red).
Figure 14: Normal model / Scenario 3 (M4) — Learned joint prior for all 30 replications Each replication is shown in a different color. The correlation structure of the regression coefficients used for learning the prior is $\rho(\beta_{0},\beta_{1})=0.3,\rho(\beta_{1},\beta_{2})=-0.2,$ and $\rho(\beta_{0},\beta_{2})=-0.3$. Independence between $\sigma$ and correlation coefficients is assumed.
@@ 5 Discussion & Outlook
In this paper, we introduce an extension of our simulation-based framework [6] for learning non-parametric joint priors. This extension builds on using normalizing flows [46] to learn flexible joint priors for the model parameters. Although we employ normalizing flows in this work, other generative models, such as flow matching [47] or diffusion models [9] are also compatible with our framework and can be readily used as a drop-in replacement.
Across four simulation studies, we demonstrated successful learning of joint priors based primarily on interpretable quantities that can be meaningfully provided by domain experts. One exception is the use of correlation information, which the current method requires to reasonably constrain the optimization problem. Future work is required to integrate in our method an elicitation technique that infers the correlation indirectly from the expert by asking about interpretable quantities. Recently, Mikkola et al [53] proposed a method that uses pairwise comparisons or ranking techniques to learn a multivariate prior density, which could serve as an interesting avenue for future research.
With the extension of our originally proposed simulation-based framework to support non-parametric prior distributions, we offer an elicitation framework, implemented in elicito [5], which can accommodate various types of elicitation methods within a single, unified structure. This facilitates the development of a standardized set of diagnostics and analysis tools, addressing a current gap in the literature [52].
@@
In addition to accommodating different types of prior distributions, another strength of the framework is that expert knowledge can be specified in various ways. This ranges from querying model parameters to observable quantities or any derived statistics, allowing the queried information to be tailored to the specific problem and the expertise of the domain expert. This design choice provides users with maximum flexibility in specifying the set of quantities to be elicited from the domain expert. However, it also places the responsibility on the user to define a valid and appropriate set of quantities to query from the expert [70]. We would argue that this flexibility is in general a desirable property of a method, as it avoids imposing specific modeling choices on the user. Thus, instead of requiring that the problem is framed in a predefined way, the method allows users to adapt it to their specific needs [66, 70, 68]. However, in practice, this flexibility is only effective if good recommendations, guidelines, and diagnostics are available to help users define an appropriate set of quantities to query from the expert [31].
@@
A significant challenge in this context is assessing whether the chosen set of statistics is sufficiently informative for the corresponding model. The goal is not necessarily to identify a unique prior that fully defines the generative model. Instead, the aim is to determine a set of equivalently plausible priors, supporting the assumption that a domain expert is unlikely to have a single “true” prior in mind [75]. This introduces at least two additional requirements for an elicitation method. First, it should include an efficient implementation that allows the training algorithm to be executed multiple times with different seeds, enabling the assessment of variability in the learned prior distributions. Second, the elicitation method should ideally offer additional diagnostics or advanced techniques, such as model selection or model averaging, to assist the user [24]. In this work, we propose a model averaging method in which the weights calculated for averaging can also serve as diagnostics, helping to identify potential outliers within the set of learned priors. Identifying these edge cases could provide a valuable starting point for evaluating the plausibility of the learned priors in collaboration with a domain expert.
@@
Other important design choices within our framework that require appropriate default settings include the specification of the discrepancy measure and the selection of the optimization approach, among other aspects. So far, in all simulation studies, we have used the maximum mean discrepancy [MMD, 38] to evaluate the discrepancy between the model-implied and expert-elicited statistics. The only exception is the correlation information, for which we employed an L2 loss. One key advantage of the MMD is its flexibility in handling a wide range of statistics, from simple point estimates to complex histogram data. However, a notable limitation of the MMD is its high (i.e., quadratic) computational cost. In contrast, the L2 loss is computationally efficient but lacks the capability to process histogram data. Despite the higher computational cost, the MMD consistently showed very good performance and robustness across all simulation studies so far. Therefore, it appears to be a solid choice as a default discrepancy measure. Future work should study the performance of different loss functions depending on different elicited statistics, with the goal to develop user guidelines for selecting an appropriate loss function.
@@
Finally, while our current methodology learns prior distributions based on expert data, it does currently not account for additional ‘meta’ uncertainty that arises during the elicitation process. This includes uncertainties introduced by the elicitation technique itself or the expert’s uncertainty regarding the ‘true’ quantity [see e.g., 24, 25, 71]. We acknowledge that these are important sources of information that should be incorporated in future developments of our method. One promising approach is to view this problem from a Bayesian perspective and treat the proposed workflow as a Bayesian inference problem. This idea was already proposed by Mikkola et al [52] and referred to as the supra-Bayesian approach. We believe this is an important and valuable direction for the next steps in our elicitation method.
@@ Acknowledgements
We thank Dmytro Perepolkin and Luna Fazio for their valuable comments and suggestions that greatly improved the manuscript.
@@
The simulation results are available on OSF at https://osf.io/xrzh6/.
@@
Supplementary material is available on GitHub at https://github.com/florence-bockting/non-parametric-prior.
@@
Code for running the simulations and plotting the results is available on GitHub at https://github.com/florence-bockting/non-parametric-prior. Simulations and most plots are based on/provided by our Python package elicito [5, v0.3.1]. To ensure the reproducibility of our results, the specific version of elicito used in this study has been archived on Zenodo and is accessible via the following DOI https://doi.org/10.5281/zenodo.15241853.
@@
Conceptualization: PB, SR, FB, Methodology: PB, FB, Software: FB, Writing (Original Draft): FB, Writing (Editing): FB, SR, PB, Writing (Review): FB, PB, Visualization: FB, Supervision: PB, Funding acquisition: PB
@@ A.1 Training time
The median training time across the 30 replications ($\pm$ sd of training time):
  • Binomial model: 6.87 minutes ($\pm$ 1.45).
  • Independent normal scenario: 10.18 minutes ($\pm$ 1.68).
  • Skewed normal scenario: 14.45 minutes ($\pm$ 2.80).
  • Correlated normal scenario: 10.10 minutes ($\pm$ 0.17).
@@ A.2 Independent normal scenario
Figure 15: Normal model / Scenario 1 (M2) — Sensitivity Analysis: Rows represent hyperparameters of each model parameter. Columns represent elicited statistics: five quantiles for $y\mid\text{gr}_{i}$ with $i=1,2,3$ and for $R^{2}$. Quantiles are depicted in different colors. In each row, the corresponding hyperparameter, is varied across the range shown on the x-axis, while all other hyperparameters are held constant at their true values, indicated by the red vertical line.
Figure 16: Normal model / Scenario 1 (M2) — Preliminary Convergence Check: Upper plot displays the absolute slope of the total loss trajectory (scaled by a factor of 100) over the last 100 epochs on the y-axis, with each replication shown along the x-axis. The replications (i.e., seeds) are arranged in ascending order based on the magnitude of the absolute slope. The best-performing model is seed 14, while the worst-performing models are seeds 6, 26, 10, and 9. The lower plot illustrates the loss trajectories of the best- and worst-performing seeds over the last 300 epochs. The loss updates for the final 100 epochs are highlighted in green (best model) and blue (worst models), with a red line segment indicating the respective slope.
Figure 17: Normal model / Scenario 1 (M2) — Visual convergence check 1: Updating trajectory across epochs for the total loss (left) and the individual (weighted) loss components (right).
Figure 18: Normal model / Scenario 1 (M2) — Visual convergence check 2: Updating trajectory across epochs for the mean (left) and standard deviation (right) of the final learned marginal priors for all 30 replications.
Figure 19: Normal model / Scenario 1 (M2) — Joint prior: Each replication is visualized in a different color.
@@ A.3 Skewed normal scenario
Figure 20: Normal model / Scenario 2 (M3) — Sensitivity Analysis: Rows represent hyperparameters of each model parameter. Columns represent elicited statistics: five quantiles for $y\mid\text{gr}_{i}$ with $i=1,2,3$ and for $R^{2}$. Quantiles are depicted in different colors. In each row, the corresponding hyperparameter, is varied across the range shown on the x-axis, while all other hyperparameters are held constant at their true values, indicated by the red vertical line.
Figure 21: Normal model / Scenario 2 (M3) — Preliminary Convergence Check for all 30 replications: The upper plot displays the absolute slope of the total loss trajectory (scaled by a factor of 100) over the last 100 epochs on the y-axis, with each replication shown along the x-axis. The replications (i.e., seeds) are arranged in ascending order based on the magnitude of the absolute slope. The best-performing model is seed 13, while the four worst-performing models are seeds 4, 10, 3, and 16. The lower plot illustrates the loss trajectories of the best- and worst-performing seeds over the last 300 epochs. The loss updates for the final 100 epochs are highlighted in green (best seed) or blue (worst seeds). The red line segment indicates the respective slope.
Figure 22: Normal model / Scenario 2 (M3) — Visual convergence check 1: Updating trajectory across epochs for the total loss (left) and the individual (weighted) loss components (right).
Figure 23: Normal model / Scenario 2 (M3) — Visual convergence check 2: Updating trajectory across epochs for the mean (left) and standard deviation (right) of the marginal priors.
Figure 24: Normal model / Scenario 2 (M3) — Visual convergence check 2: Joint prior distribution. Each replication is visualized in a different color.
@@ A.4 Correlated normal scenario
Figure 25: Normal model / Scenario 3 (M4) — Sensitivity Analysis: Rows represent hyperparameters of each model parameter. Columns represent elicited statistics: five quantiles for $y\mid\text{gr}_{i}$ with $i=1,2,3$ and for $R^{2}$. Quantiles are depicted in different colors. In each row, the corresponding hyperparameter, is varied across the range shown on the x-axis, while all other hyperparameters are held constant at their true values, indicated by the red vertical line.
Figure 26: Normal model / Scenario 3 (M4) — Preliminary Convergence Check for all 30 replications: The upper plot displays the absolute slope of the total loss trajectory (scaled by a factor of 100) over the last 100 epochs on the y-axis, with each replication shown along the x-axis. The replications (i.e., seeds) are arranged in ascending order based on the magnitude of the absolute slope. The best-performing model is seed 6, while the worst-performing models are seeds 11, 0, 13, and 20. The lower plot illustrates the loss trajectories of the best- and worst-performing seeds over the last 200 epochs. The loss updates for the final 100 epochs are highlighted in green for the best model and in blue for the worst models, with a red line segment indicating the respective slope.
Figure 27: Normal model / Scenario 3 (M4) — Visual convergence check 1: Updating trajectory across epochs for the total loss (left) and the individual (weighted) loss components (right).
Figure 28: Normal model / Scenario 3 (M4) — Visual convergence check 1: Updating trajectory across epochs for the mean (left) and standard deviation (right) of the marginal priors.
Figure 29: Normal model / Scenario 3 (M4) — Visual convergence check 2: Joint prior distribution. Each replication is visualized in a different color.